Consider the population consisting of values

3, 5, and 7.


List all the possible samples of size 2 that


can be drawn from the population with


replacement in the table below.


b. Find the mean of the sampling distribution


of means.


c. Find the variance of the sampling


distribution of the means


d. Find the standard deviation of the means


1
Expert's answer
2022-03-29T07:50:52-0400

"\mu(3,5)=(3+5)/2=4"

"\mu(3,7)=(3+7)/2=5"

"\mu(5,7)=(5+7)/2=6"

"\mu(3,3)=(3+3)/2=3"

"\mu(5,5)=(5+5)/2=5"

"\mu(7,7)=(7+7)/2=7"

b. f(4)=f(6)=f(3)=f(7)=1/6

f(5)=2/6=1/3

"E(x)=\sum f\mu=1/6(4+6+3+7)+2/6\times5=5"

"\sigma^2=\sum f\mu^2-(\sum f\mu)^2=1/6(16+36+9+49)+2/6\times25-25=1.67"

"\sigma=\sqrt{\sigma^2}=\sqrt{1.67}=1.29"